Date: Thu, 27 Aug 1998 11:07:46 -0700 (PDT)
To: 2mass
Subject: IPAC 2MASS WG Meeting #162 Minutes
Cc: chas, stiening, bgreen
Mime-Version: 1.0
Content-Transfer-Encoding: quoted-printable
Content-MD5: upynQB1tQ5c+opZG4KnLYA==

           IPAC 2MASS Working Group Meeting #162 Minutes

Attendees: R. Cutri, S. Van Dyk, R. Beck, T. Evans, R. Hurt, J. White,
           J. Fowler, G. Kopan, H. McCallon, S. Wheelock, R. Tam, 
           D. Kirkpatrick, W. Wheaton, T. Jarrett


1.) Project Update
2.) 2MAPPS 2.1 Delivery
3.) 2MASS Sampler
4.) QA Update


     The meeting began with distribution of rather tasty materials brought by
D. Kirkpatrick from his most recent trip to the Big Island, Hawaii.

1.) Project Update

     The southern telescope was closed, due to a dome drive failure, which is
being repaired.  Guenter Riegler was visiting this week on WIRE business, but
R. Cutri was to provide him with a 2MASS update as well.

2.) 2MAPPS 2.1 Delivery

     R. Cutri inquired of W. Wheaton on new news on the PSF generation front.
W. Wheaton responded that he had no new PSFs, but was taking inventory of the
current PSFs and rerunning the seeing history.  R. Cutri stressed that 
finalizing the southern PSFs was a high-priority action item.  W. Wheaton will
intends to have a whole new set of PSFs by midday Friday.  He intends to make
determination of the saturation thresholds on the south arrays the next order
of business.

[CLARIFICATION:  W. Wheaton will only be filling out the table of PSFs
for bad seeing (i.e., seeing "shape" parameter sh > 1.01 or so), 
not delivering a whole new set.  The PSFs already delivered are OK, 
and they cover the seeing frequency histograms for the
south up to about the 90-95 percentile level. That is, 5-10% of
the data have seeing worse than that covered by the PSF table. 

Because bad seeing is relatively rare,  a lot of nights in the south
must be inventoried to be sure of having enough data to make PSFs 
that cover the tail of the histogram.]

     R. Cutri provided an update on the photometric bias fix.  G. Kopan has a
preliminary version of code based on analyses of several nights.  For a block of
scans (each set of cal scans and sci scans) the routine generates a bias profile
over 10 cross-scan bins, performing linear interpolation within the profile 
bins.  From this profile it then generates correction tables.  S. Wheelock will
provide code to apply the corrections to the point source files.  G. Kopan hopes
to have the code ready by Wednesday.  An off-chance exists that there may need 
to be a new MAPCOR that rederives the psf-fit vs. aperture photometry
normalization after the cross-scan variation correction has been made.  R. Cutri
indicated to T. Evans that she contemplate on this, but not take any immediate 
action.  The normalization and calibration issues that result from the 
correction will will be analyzed when the prototype correction code has been 

     It is possible that the bias correction could be established based on a
dependence of the bias on PSF ID's.  This could be determined only for dense
scans, since, as G. Kopan pointed out, some cal scan blocks have too small an
amount of information to acquire enough statistics.  G. Kopan will look into
the PSF ID relation for some dense scans.  

3.) 2MASS Sampler

     The Sampler night 971116n is being rerun, due to a MAPCOR error.  T. Evans
is to determine the magnitude of this error, as it affects  the data that have 
been ops-processed since July 16. The glint artifact parameters are "out of 
whack," based on a new namelist which was in the wrong directory.  Havoc is not 
being wreaked on the entire glint search, but H and Ks secondary glints may be 
lost, due to the current magnitude threshold.  The problem seems to be only for
the north.  The correction will have to be applied to the database.

     R. Beck will issue a memo on the rules for redelivery and namelist
locations.  The src deliveries should be made to /2massc/del/src and the  
namelist files to the datan and datas directories, on rodan.  R. Beck is working
on a change log webpage, that he will maintain.  He stated that he would prefer
no changes in namelists.  But team members will be alerted when major changes
in deliveries are made.

     One potentially disturbing aspect of the most recent run of 971116n showed
that the number of sources in the .bfpts file (for, say, one scan) was 
different than for the previous run.  Why this is so is not clear and must be
hunted down.  One test to locate these "new" sources would be to run smerge on
both the old and new .bfpts files.

     Because of these problems with the most recent run and the need for yet
another re-run, the Sampler has now slipped into October, rather than the
intended mid-September date.

     T. Jarrett is generating the extended sources catalog.  The catalog will
consist of a >300-column table with various parameters.  T. Jarrett will provide
to T. Evans the criteria with thresholds on these parameters for the auxiliary
extended source classification table, which will consist of about 10 columns.
There will be both a short and long form of this catalog.  The first "view" will
be the short form, with option to go to a second, longer "view".

     R. Cutri gave as "homework" to *everyone* on the team to test out the IRSA
interface, by trying out queries in CatScan and utilizing the Survey Visualizer.
Any suggestions or comments on how to improve this interface should be directed
to J. Mazzarella and J. Good, being sure to "cc" R. Cutri and T. Handley on any

4.) QA Update

     R. Hurt will be working soon on the scan-by-scan QA webpage.